A Restaurant manager decides to lower the price of a featured sandwich from Rs. 500 

to Rs. 400, and she finds that sales during the week increase from 440 to 680 

sandwiches. Is demand elastic or Inelastic?


1
Expert's answer
2021-05-09T14:32:03-0400

Elasticity: It is used for measuring the change in the behavior of buyers and sellers due to a change in price for a good or service.

Initial price= 500

New price=400

Initial quantity= 440

New quantity= 680

"PED=\frac{change \space in\space quantity}{change\space in \space price}\times\frac{P_1+P_2}{Q_1+Q2}"


"PED=\frac{680-440}{400-500}\times \frac{500+400}{440+680}"


"PED=\frac{240}{-100}\times \frac{900}{1120}"


"PED=-2.4\times 0.8036"


"PED=-1.93"


The demand for Sandwiches is elastic because it changes drastically when its price decreases.